# Artificial Intelligence in Image-based Cardiovascular Disease Analysis: A Comprehensive Survey and Future Outlook

### v3

@@ Abstract
Recent advancements in Artificial Intelligence (AI) have significantly influenced the field of Cardiovascular Disease (CVD) analysis, particularly in image-based diagnostics. Our paper presents an extensive review of AI applications in image-based CVD analysis, offering insights into its current state and future potential. We systematically categorize the literature based on the primary anatomical structures related to CVD, dividing them into non-vessel structures (such as ventricles and atria) and vessel structures (including the aorta and coronary arteries). This categorization provides a structured approach to explore various imaging modalities like Magnetic Resonance Imaging (MRI), which are commonly used in CVD research. Our review encompasses these modalities, giving a broad perspective on the diverse imaging techniques integrated with AI for CVD analysis. Additionally, we compile a list of publicly accessible cardiac image datasets and code repositories, intending to support research reproducibility and facilitate data and algorithm sharing within the community. We conclude with an examination of the challenges and limitations inherent in current AI-based CVD analysis methods and suggest directions for future research to overcome these hurdles.
Index Terms:
AI, Cardiac Image, Cardiovascular Diseases, Deep Learning, Medical Image Analysis, Survey.
@@ 1 Introduction
Cardiovascular diseases (CVDs) are the leading cause of mortality worldwide, accounting for an estimated 17.9 million deaths annually, as per the World Health Organization (WHO). In the United States, CVDs, encompassing heart disease and stroke, rank as the primary cause of death. Heart disease alone was responsible for approximately 697,000 deaths in 2020, according to the Centers for Disease Control and Prevention (CDC). These statistics underscore the profound impact of CVDs on global public health and underscore the critical need for enhanced medical research and healthcare strategies in this domain. Consequently, timely and accurate diagnosis, risk assessment, and treatment planning are essential to mitigate their impact and improve patient outcomes [1, 2].
@@
Cardiac imaging encompasses a variety of medical imaging modalities designed to visualize and assess the heart’s structure and function [3]. These modalities offer detailed insights into the heart’s chambers, blood vessels, and associated tissues, serving multiple critical functions in healthcare. (i) Cardiac imaging aids in identifying biomarkers linked to the risk, progression, and treatment response of heart diseases [4]. Indicators such as carotid artery wall thickness can act as a biomarker for atherosclerosis and predict future cardiovascular events. It can also reveal changes in heart tissue or blood flow that serve as biomarkers for various heart conditions [5], enhancing disease understanding and guiding treatment decisions. (ii) Cardiac imaging is integral to both the planning and monitoring of treatments for various cardiovascular diseases, including coronary artery disease, heart failure, and valvular heart disease. Techniques like angiography are instrumental in visualizing blood flow in coronary arteries and detecting obstructions, while echocardiography evaluates heart function and detects structural or motion abnormalities. These imaging methods enable physicians to devise optimal treatment strategies and adjust them over time based on patient progress. (iii) In the realm of clinical trials, cardiac imaging plays a pivotal role in evaluating the safety and efficacy of new medical therapies. It allows for real-time visualization of the heart’s internal structures and functions and can track morphological or functional changes induced by new drugs or treatments. This capability is essential for determining the effectiveness and safety of treatments before their widespread adoption.
@@
The incorporation of Artificial Intelligence (AI) in medical imaging has significantly transformed the analysis of image-based CVDs [6, 7]. AI, particularly deep learning methods [8, 9, 10], has markedly improved the precision, efficiency, and objectivity of interpreting cardiovascular images [11, 12]. The integration of AI with various cardiovascular imaging techniques, including MRI, CT, X-ray, and ultrasound, has enabled more comprehensive and dynamic evaluations of cardiovascular structures and functions [13, 14]. Recent AI advancements have facilitated breakthroughs in cardiovascular imaging tasks, including segmentation, disease classification, risk prediction, and clinical decision-making support for treatment planning [11, 15]. These developments highlight AI’s crucial role in enhancing the fight against cardiovascular diseases.
@@
In this comprehensive review, we delve into the recent advancements and trends in Artificial Intelligence (AI) applied to the image-based analysis of cardiovascular diseases (CVDs), as illustrated in Figure 1. Our approach is threefold: (i) Categorization Based on Anatomical Structures: We systematically organize an extensive collection of CVD research according to the primary anatomical structures impacted and their functions. This categorization creates two principal groups: non-vessel structures such as atria and ventricles, and vessel structures, including the aorta and coronary arteries (See Figure 1). This classification is based on similarities in both the anatomical features and the analysis techniques employed. For example, vessel structure analyses often involve methods like vessel tracing and stenosis degree estimation. (ii) Integration with Other Data Types: Beyond traditional medical imaging, our review includes studies that integrate imaging with additional data types, like genomics, for a more holistic analysis of CVDs [16]. This expansive coverage, which includes a variety of cardiac imaging structural analysis tasks—such as chamber segmentation, coronary artery segmentation, and coronary artery branch labeling—sets our review apart. Additionally, we explore functional simulation methods, such as fractional flow reserve (FFR), a vital diagnostic measure in cardiology for assessing the severity of coronary artery stenosis. (iii) Large-Scale Population-Based Studies: The review expands to encompass large-scale studies using AI for image-based analysis of CVDs. The merger of imaging genetics with AI in cardiovascular disease research represents a significant leap forward. It provides researchers with the tools to decode the genetic foundations of these diseases, significantly improving diagnosis, treatment strategies, and prognostic evaluations [16].
@@
There exists a range of survey papers that have provided overviews of AI applications in cardiovascular imaging tasks, such as segmentation [17, 18, 19], or that have focused on specific diseases [20], types of CVD like Congenital heart disease (CHD) [21], or singular medical image modalities [22]. However, these surveys have typically lacked a comprehensive approach that encompasses both structural and functional aspects of cardiovascular analysis. Notably, critical techniques like coronary flow analysis [23] are often omitted. In our survey, we address this gap by discussing cardiac imaging modalities along with their unique characteristics, advantages, and the diseases they are most relevant to. Our approach covers both the anatomical aspects (such as chamber volume and coronary artery morphology) and functional aspects (including cardiac motion and blood flow) of cardiovascular imaging. Additionally, we provide a summary of publicly available datasets and code repositories to enhance research reproducibility, foster collaboration, and accelerate scientific advancements in this field. Furthermore, our survey engages in insightful discussions about the current challenges in AI-based cardiovascular imaging and outlines potential directions for future research in this rapidly evolving domain.
@@
Figure 1: This figure illustrates our comprehensive approach to exploring cardiovascular diseases (CVDs). We categorize these diseases according to the primary anatomical structures they impact and their respective functions. Our analysis also encompasses a variety of medical imaging modalities employed in the diagnosis and study of these diseases. Furthermore, we highlight the significant role of Artificial Intelligence (AI) in enhancing image-based cardiovascular disease analysis, showcasing the integration of advanced AI techniques with traditional imaging methods. The example categorizes CVDs by highlighting the key anatomical structures involved and their functions, also showcasing the data modalities and AI-based methods used for analysis. Top: It includes examples of non-vessel anatomical structures, such as (a) atria [20], (b) ventricles [24] and (c) Myocardium [25]. Bottom: It presents vessel structures, including the (d) aorta [26] and (e) coronary arteries [27], [28].
@@
  • A significant aspect of our work is the integration of cardiac imaging with large-scale, population-based studies through AI technologies. We explore how AI can effectively analyze image-based CVD data in combination with genetic, lifestyle, and environmental health factors. This multidisciplinary approach results in a nuanced understanding of CVD, considering the interplay between medical imaging data and broader genetic and environmental factors. Such comprehensive analysis is crucial for developing more personalized and effective strategies for CVD prevention and management.
  • Our survey also sheds light on potential future research directions, particularly emphasizing the need for integrating diverse data types. This includes combining environmental and genetic data to develop sophisticated AI models for addressing complex clinical challenges in heart disease. By leveraging such multi-scale data, AI models can significantly improve our understanding of heart diseases, aiding in early detection, accurate diagnosis, and the formulation of effective treatment plans. This suggests that future advancements in AI will likely concentrate on leveraging these diverse data modalities to address intricate clinical questions in cardiology, aiming to enhance patient outcomes significantly.
The remaining content of this paper is organized as follows. After introducing cardiac imaging modalities as the background in Section 2, we describe the heart structures (non-vessel and vessel), the common diseases, and the related AI works in Section 3. We also review the population-based cardiovascular imaging analysis works in Section 4. In Sections 5, we summarize the public datasets and code. We discuss the limitations of the current AI-based methods and the future works in Section 6. Finally, we conclude the paper in Section 7.
@@ 2 Background - Cardiac Imaging
In this section, we delve into the commonly employed modalities in cardiac imaging (examples are shown in Figure 2). Depending on the specific clinical requirements and the information needed by healthcare professionals, these imaging techniques are either used in combination or selectively chosen to provide the most accurate and comprehensive assessment of cardiac health.
Figure 2: Examples of Cardiac Data Modalities. Top (Non-vessel): Cardiac MRI [29], Cine MRI and Tagged Cine MRI [30], Nuclear Cardiology [31], Cardiac Ultrasound [32]. Bottom (Vessel): Cardiac CT (Aorta [33], Coronary [34, 35]), Cardiac X-ray [36], Intravascular Ultrasound (IVUS) [37].
@@ 2.1 Electrocardiogram (ECG)
An electrocardiogram (ECG) [38] is a critical medical test that captures the heart’s electrical activity [39]. It graphically represents this activity over time, with the x-axis denoting time and the y-axis the amplitude of the electrical signal. The ECG can be perceived as a form of medical imaging, mapping the heart’s electrical impulses through wave patterns on a graph [40]. To obtain an ECG, electrodes are strategically placed on the patient’s skin to record electrical impulses generated during the heart’s contraction and relaxation cycles. Each heartbeat begins with an electrical impulse originating from the heart’s sinoatrial (SA) node. This impulse travels through the heart muscle, prompting it to contract and pump blood. The ECG machine detects and amplifies these impulses, transforming them into waveforms that offer insights into heart function.
ECG is an indispensable diagnostic tool in cardiovascular medicine, offering clinicians immediate insight into the heart’s electrical function [41]. It is essential for diagnosing a variety of heart-related conditions, including: (i) Heart Rhythm Disorders: An ECG is effective in identifying abnormal heart rhythms or arrhythmias, such as atrial fibrillation or ventricular tachycardia. (ii) Heart Attacks: The ECG is instrumental in recognizing patterns indicative of myocardial ischemia (reduced blood flow) or infarction (tissue death due to lack of oxygen), which are key markers of heart attacks. (iii) Structural Heart Abnormalities: An ECG can reveal structural heart changes that suggest conditions like cardiomyopathy or hypertrophy, where the heart muscle is enlarged or altered [42].
2.2 Cardiac X-ray Imaging
Cardiac X-ray imaging (can be static images and/or real-time fluoroscopic imaging), often employed in coronary angiography, is a diagnostic technique that uses X-rays to produce two-dimensional images of the heart and blood vessels [43]. This method is particularly useful in emergency settings for initial assessments due to its quick execution and relatively low radiation exposure, making it safer for a wide range of patients [44]. It is excellent for identifying coronary artery disease (CAD) by directly visualizing the coronary arteries and informing treatment strategies like angioplasty or stenting to restore heart muscle blood flow [45].
Digital Subtraction Angiography (DSA): This advanced technique in cardiac X-ray imaging enhances image clarity by digitally removing background structures, focusing solely on blood vessels. The use of contrast dye in conjunction with rapid image acquisition results in high-resolution images that are crucial for identifying vascular abnormalities [46].
Interventional Coronary Angiography (ICA): ICA is a specialized application of Cardiac X-ray imaging used for interventional procedures such as percutaneous coronary intervention (PCI). During PCI, a catheter is inserted to treat blockages in coronary arteries, often employing techniques like angioplasty and stenting. The X-ray system plays a crucial role in guiding the catheter’s placement and monitoring the progress of the intervention [47].
@@ 2.3 Cardiac Computed tomography (CT)
Cardiac Computed Tomography (CT), utilizing advanced X-ray technology, provides detailed cross-sectional, three-dimensional images of the heart and surrounding vessels. Compared to Cardiac X-ray imaging, Cardiac CT involves higher levels of radiation exposure but offers more detailed and comprehensive views. This imaging method is particularly valuable for assessing cardiac anatomy, with a primary focus on the coronary arteries. It plays a crucial role in identifying blockages or narrowing within these arteries, which are key indicators of coronary artery disease [48]. In the following sections, we will explore various Cardiac CT modalities and their specific applications in cardiac imaging [49].
Coronary CT Angiography (CTA): Non-contrast CT imaging utilizes tissue density differences to create images, allowing distinction between soft tissues, calcium, fat, and air. This is useful for estimating calcium presence in coronary arteries. In contrast-enhanced coronary CTA, acquired post-contrast agent injection, the imaging provides detailed views of cardiac chambers, vessels, and coronaries, effectively detecting non-calcified coronary plaques [50]. It’s widely employed for diagnosing coronary artery disease (CAD), coronary anomalies, and pre-surgical evaluations for coronary bypass. Additionally, it assesses stent patency post-implantation [51].
Calcium-Scoring Heart Scan: This scan, also known as coronary artery calcium (CAC) scoring, is a specialized X-ray test that measures calcifications in coronary arteries, indicative of coronary artery disease (CAD). These calcifications, even before symptom onset, signal potential heart-related events. The derived calcium score helps in risk assessment for CAD, with higher scores suggesting increased risk of heart attacks [52].
Functional Cardiovascular CT: This CT form is essential for non-invasive evaluation of heart and vascular structures. Known for producing detailed body cross-sections, it offers insights into both structural and functional aspects of the cardiovascular system. Functional Cardiovascular CT is instrumental in assessing myocardial perfusion and ventricular function, useful in evaluating ischemic heart disease, cardiomyopathies, and heart failure [53].
Cardiac Dual-Energy CT (DECT): DECT applies dual-energy imaging principles to the heart and vasculature, utilizing two X-ray energy spectra to yield detailed information on cardiac tissues and pathologies. It enhances plaque characterization in arteries and differentiates tissue types, aiding in myocardial infarction diagnosis and tumor identification [54].

### v4

@@ Abstract
Recent advancements in Artificial Intelligence (AI) have significantly influenced the field of Cardiovascular Disease (CVD) analysis, particularly in image-based diagnostics. Our paper presents an extensive review of AI applications in image-based CVD analysis, offering insights into its current state and future potential. We systematically categorize the literature based on the primary anatomical structures related to CVD, dividing them into non-vessel structures (such as ventricles and atria) and vessel structures (including the aorta and coronary arteries). This categorization provides a structured approach to explore various imaging modalities like Computed tomography (CT) and Magnetic Resonance Imaging (MRI), which are commonly used in CVD research. Our review encompasses these modalities, giving a broad perspective on the diverse imaging techniques integrated with AI for CVD analysis. We conclude with an examination of the challenges and limitations inherent in current AI-based CVD analysis methods and suggest directions for future research to overcome these hurdles.
keywords:
Artificial Intelligence, Cardiac Image, Cardiovascular Diseases, Deep Learning, Survey.
doi:
10.1146/…
journal: The Annual Review of Biomedical Data Science.
@@ 1 Introduction
Cardiovascular disease (CVD) is the leading cause of mortality worldwide, accounting for an estimated 17.9 million deaths annually, as per the World Health Organization (WHO). CVD, encompassing heart disease and stroke, is the primary cause of death in the United States. Heart disease alone was responsible for approximately 697,000 deaths in 2020, according to the Centers for Disease Control and Prevention (CDC). These statistics underscore the profound impact of CVDs on global public health and underscore the critical need for enhanced medical research and healthcare strategies in this domain. Consequently, timely and accurate diagnosis, risk assessment, and treatment planning are essential to mitigate their impact and improve patient outcomes [update2017heart, sun2023social].
@@
Cardiac imaging encompasses a wide range of modalities used to visualize and evaluate the heart’s structure and function [li2023multi]. These methods provide essential information about heart’s chambers, blood vessels, and associated tissues, supporting several key roles in clinical care and research. (i) Cardiac imaging aids in identifying biomarkers linked to the risk, progression, and treatment response of heart diseases [vasan2006biomarkers]. Indicators such as carotid artery wall thickness can act as a biomarker for atherosclerosis and predict future cardiovascular events. It can also reveal changes in heart tissue or blood flow that serve as biomarkers for various heart conditions [michelhaugh2023using], enhancing disease understanding and guiding treatment decisions. (ii) Cardiac imaging is integral to both the planning and monitoring of treatments for various cardiovascular diseases, including coronary artery disease, heart failure, and valvular heart disease. Techniques like angiography are instrumental in visualizing blood flow in coronary arteries and detecting obstructions, while echocardiography evaluates heart function and detects structural or motion abnormalities. These imaging methods enable physicians to devise optimal treatment strategies and adjust them over time based on patient progress. (iii) In the realm of clinical trials, cardiac imaging plays a pivotal role in evaluating the safety and efficacy of new medical therapies. It allows for real-time visualization of the heart’s internal structures and functions and can track morphological or functional changes induced by new drugs or treatments. This capability is essential for determining the effectiveness and safety of treatments before their widespread adoption.
@@
The incorporation of Artificial Intelligence (AI) in medical imaging has significantly transformed the analysis of CVDs [bizopoulos2018deep, krittanawong2023deep]. AI, particularly deep learning methods [lecun2015deep, mascarenhas2021comparison, zhou2021review], has markedly improved the precision, efficiency, and objectivity of interpreting cardiovascular images [shen2017deep, yang2023application]. The integration of AI with various cardiovascular imaging techniques, including MRI, CT, X-ray, and ultrasound, has enabled more comprehensive and dynamic evaluations of cardiovascular structures and functions [he2023blinded, litjens2019state]. Recent AI advancements have facilitated breakthroughs in cardiovascular imaging tasks, including segmentation, disease classification, risk prediction, and clinical decision-making support for treatment planning [shen2017deep, shin2016deep]. These developments highlight AI’s crucial role in enhancing the fight against cardiovascular diseases.
@@
In this comprehensive review, we delve into the recent advancements and trends in Artificial Intelligence (AI) applied to the image-based analysis of cardiovascular diseases (CVDs), as illustrated in Figure 1. Our approach is threefold: (i) Categorization Based on Anatomical Structures: We systematically organize an extensive collection of CVD research according to the primary anatomical structures impacted and their functions. This categorization creates two principal groups: non-vessel structures such as atria and ventricles, and vessel structures, including the aorta and coronary arteries (See Figure 1). This classification is based on similarities in both the anatomical features and the analysis techniques employed. For example, vessel structure analyses often involve methods like vessel tracing and stenosis degree estimation. (ii) Integration with Other Data Types: Beyond traditional medical imaging, our review includes studies that integrate imaging with additional data types, like genomics, for a more holistic analysis of CVDs [zhao2023heart]. This expansive coverage, which includes a variety of cardiac imaging structural analysis tasks—such as chamber segmentation, coronary artery segmentation, and coronary artery branch labeling—sets our review apart. Additionally, we explore functional simulation methods, such as fractional flow reserve (FFR), a vital diagnostic measure in cardiology for assessing the severity of coronary artery stenosis. (iii) Large-Scale Population-Based Studies: The review expands to encompass large-scale studies using AI for image-based analysis of CVDs. The merger of imaging genetics with AI in cardiovascular disease research represents a significant leap forward. It provides researchers with the tools to decode the genetic foundations of these diseases, significantly improving diagnosis, treatment strategies, and prognostic evaluations [zhao2023heart].
@@
There exists a range of survey papers that have provided overviews of AI applications in cardiovascular imaging tasks, such as segmentation [chen2020deep, fu2021review], or that have focused on specific diseases [li2022medical], types of CVD like congenital heart disease (CHD) [jone2022artificial], or singular medical image modalities [jafari2023automated]. However, these surveys have typically lacked a comprehensive approach that encompasses both structural and functional aspects of cardiovascular analysis. Notably, critical techniques like coronary flow analysis [berry2014fractional] are often omitted. In our survey, we address this gap by discussing cardiac imaging modalities along with their unique characteristics, advantages, and the diseases they are most relevant to. Our approach covers both the anatomical aspects (such as chamber volume and coronary artery morphology) and functional aspects (including cardiac motion and blood flow) of cardiovascular imaging. Furthermore, our survey engages in insightful discussions about the current challenges in AI-based cardiovascular imaging and outlines potential directions for future research in this rapidly evolving domain.
@@
Figure 1: This overview figure illustrates our comprehensive approach to exploring cardiovascular diseases (CVDs). We categorize these diseases according to the primary anatomical structures they impact and their respective functions. Our analysis also encompasses a variety of medical imaging modalities employed in the diagnosis and study of these diseases. The example categorizes CVDs by highlighting the key anatomical structures involved and their functions. Top: It includes examples of non-vessel anatomical structures, such as (a) atria [li2022medical], (b) ventricles [leclerc2019deep] and (c) Myocardium [ye2021deeptag]. Bottom: It presents vessel structures, including the (d) aorta [pirola20194] and (e) coronary arteries [serrano2023coronary], [kolli2018image].
@@
  • Another key contribution of our work lies at the intersection of cardiac imaging and population studies through AI, where we systematically integrate image-based CVD data with multi-scale determinants of health, including genetic predispositions, lifestyle factors, etc. This multidisciplinary paradigm enables a holistic understanding of CVD pathogenesis by deciphering the interactions between imaging phenotypes and broader biological/social contexts. Such integration is not only critical for personalizing prevention strategies but also reveals unmet needs in current research: future advances must prioritize developing AI architectures capable of synthesizing these heterogeneous data modalities to address clinically complex scenarios.
@@ 2 Background - Cardiac Imaging
In this section, we delve into the commonly employed modalities in cardiac imaging. Depending on the specific clinical requirements and the information needed by healthcare professionals, these imaging techniques are either used in combination or selectively chosen to provide the most accurate and comprehensive assessment of cardiac health.
@@ 2.1 Electrocardiogram (ECG)
An electrocardiogram (ECG) [shenoy2024novel, christensen2024vision] is a critical medical test that captures the heart’s electrical activity [berkaya2018survey]. It graphically represents this activity over time, with the x-axis denoting time and the y-axis the amplitude of the electrical signal. The ECG can be perceived as a form of medical imaging, mapping the heart’s electrical impulses through wave patterns on a graph [li2020survey]. To obtain an ECG, electrodes are strategically placed on the patient’s skin to record electrical impulses generated during the heart’s contraction and relaxation cycles. Each heartbeat begins with an electrical impulse originating from the heart’s sinoatrial (SA) node. This impulse travels through the heart muscle, prompting it to contract and pump blood. The ECG machine detects and amplifies these impulses, transforming them into waveforms that offer insights into heart function.
2.2 Cardiac X-ray
Cardiac X-ray imaging (can be static images and/or real-time fluoroscopic imaging), often employed in coronary angiography, is a diagnostic technique that uses X-rays to produce two-dimensional images of the heart and blood vessels [matsuoka2022deep]. This method is particularly useful in emergency settings for initial assessments due to its quick execution and relatively low radiation exposure, making it safer for a wide range of patients [ccimen2016reconstruction]. It is excellent for identifying coronary artery disease (CAD) by directly visualizing the coronary arteries and informing treatment strategies like angioplasty or stenting to restore heart muscle blood flow [yu2018coronary].
2.2.1 Digital Subtraction Angiography (DSA)
This advanced technique in cardiac X-ray imaging enhances image clarity by digitally removing background structures, focusing solely on blood vessels. The use of contrast dye in conjunction with rapid image acquisition results in high-resolution images that are crucial for identifying vascular abnormalities [crabb2023deep].
2.2.2 Interventional Coronary Angiography (ICA)
ICA is a specialized application of Cardiac X-ray imaging used for interventional procedures such as percutaneous coronary intervention (PCI). During PCI, a catheter is inserted to treat blockages in coronary arteries, often employing techniques like angioplasty and stenting. The X-ray system plays a crucial role in guiding the catheter’s placement and monitoring the progress of the intervention [niimi2022machine].
@@ 2.3 Cardiac Computed tomography (CT)
CT, utilizing advanced X-ray technology, provides detailed cross-sectional, three-dimensional images of the heart and surrounding vessels. Compared to Cardiac X-ray imaging, Cardiac CT involves higher levels of radiation exposure but offers more detailed and comprehensive views. This imaging method is particularly valuable for assessing cardiac anatomy, with a primary focus on the coronary arteries. It plays a crucial role in identifying blockages or narrowing within these arteries, which are key indicators of coronary artery disease [garg2023role]. In the following sections, we will explore various Cardiac CT modalities and their specific applications in cardiac imaging [gu2021cyclegan].
2.3.1 Coronary CT Angiography (CTA)
Non-contrast CT imaging utilizes tissue density differences to create images, allowing distinction between soft tissues, calcium, fat, and air. This is useful for estimating calcium presence in coronary arteries. In contrast-enhanced coronary CTA, acquired post-contrast agent injection, the imaging provides detailed views of cardiac chambers, vessels, and coronaries, effectively detecting non-calcified coronary plaques [lin2022deep]. It’s widely employed for diagnosing coronary artery disease (CAD), coronary anomalies, and pre-surgical evaluations for coronary bypass. Additionally, it assesses stent patency post-implantation [xu2023coronary].
2.3.2 Calcium-Scoring Heart Scan
This scan, also known as coronary artery calcium (CAC) scoring, is a specialized X-ray test that measures calcifications in coronary arteries, indicative of coronary artery disease (CAD). These calcifications, even before symptom onset, signal potential heart-related events. The derived calcium score helps in risk assessment for CAD, with higher scores suggesting increased risk of heart attacks [eng2021automated].
2.3.3 Functional Cardiovascular CT
This CT form is essential for non-invasive evaluation of heart and vascular structures. Known for producing detailed body cross-sections, it offers insights into both structural and functional aspects of the cardiovascular system. Functional Cardiovascular CT is instrumental in assessing myocardial perfusion and ventricular function, useful in evaluating ischemic heart disease, cardiomyopathies, and heart failure [peper2020functional].
2.3.4 Cardiac Dual-Energy CT (DECT)
DECT applies dual-energy imaging principles to the heart and vasculature, utilizing two X-ray energy spectra to yield detailed information on cardiac tissues and pathologies. It enhances plaque characterization in arteries and differentiates tissue types, aiding in myocardial infarction diagnosis and tumor identification [bruns2020deep].